6 coins are tossed simultaneously. Find the probability of getting
1) no heads
2) 3 heads

Grade:Upto college level

1 Answers

SHAIK AASIF AHAMED
9 years ago
Hello student,
As 6 coins are tossed simultaneously
sample space=26=64
For no heads all must be tails it is possible only for 1 case
P(no heads)=1/64
P(3 heads)=6c3/64=20/64
Thanks and Regards
Shaik Aasif
askIItians faculty